OpenStudy (anonymous):

two ships leave a port at 9AM. One travels at a bearing of N 53 degrees W at 12 miles per hour and the other travels at a bearing of S 67 degrees W at 16 miles per hour. Approximate hwo far apart they are at noon that day

OpenStudy (phi):

 In three hours the nw boat went 36 miles, and the sw boat went 48 miles. The angle between their two courses is 60. The easiest way to find the distance between them is to use the law of cosines: \[ c^2= a^2 + b^2 -2a \cdot c \cdot cos(C) \]